OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (1, –7); Slope -1/2

OpenStudy (twerk):

idk

OpenStudy (anonymous):

a y + 7 =2/3 (x + 1)

OpenStudy (anonymous):

y –7 =-2/3 (x + 1)

OpenStudy (anonymous):

y –7 =2/3 (x – 1)

OpenStudy (anonymous):

y + 7 =-2/3 (x – 1)

OpenStudy (anonymous):

y+7 =-1/2(x-1) y = -1/2x-13/2

OpenStudy (anonymous):

@ganeshie8 @kelliegirl33

OpenStudy (anonymous):

y - y1 = m(x - x1) <-- point slope form (1,-7) x1 = 1 and y1 = -7 m = -1/2 now just sub y - (-7) = -1/2(x - 1) y + 7 = -1/2(x - 1) ---your answer did you make a mistake when writing the problem ? Is the slope -1/2 or is it 2/3 ?

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

y - y1 = m(x - x1) m = 2/3 (1,-7) x1 = 1 and y1 = -7 now just sub y - (-7) = 2/3(x - 1) y + 7 = 2/3(x - 1) <== your answer